OpenStudy (anonymous):

Please can You help! A person is selected at random from a population that has the following characteristics are as follow: 30% of the people are women; the others are men (70%) 9% of the women are alcoholic 2% of the men are alcoholic i-)Find the probability that the selected person is a non-alcoholic, given that the person is a women. ii-)Find the probability that the selected person is a men who is alcoholic.

OpenStudy (anonymous):

Show your work first so that others can know where to offer a help from and also avoid repeating what you have already known.

OpenStudy (anonymous):

P(non-alcoholic)=1-p%w

OpenStudy (anonymous):

P(non-alcoholic)=1- (.3*.91)

OpenStudy (anonymous):

Pw(non-alcoholic)=1-(.3*.91)..?

OpenStudy (anonymous):

Given that the person is a women, what is the probability that she is alcoholic? (The answer lies inside the problem.)

OpenStudy (anonymous):

P(W-alcoholic)= .30*.09=2.7% so i) p(W-nonalcoholic)=30%-2.7%=27.3%

OpenStudy (anonymous):

No, not that way. Stare at this sentence : '9% of the women are alcoholic'. Do you know what it means?

OpenStudy (anonymous):

no explain?9%

OpenStudy (anonymous):

Well it means given that a person is a woman, the chance that she is alcoholic is 9%. Think about it.

OpenStudy (anonymous):

i)Therefore the chance the remaining women chance P(w non alcoholic) is 100%-9%=91%

OpenStudy (anonymous):

great, how about (ii)?

OpenStudy (anonymous):

correction the women population 30%-9%=21%

OpenStudy (anonymous):

what is the 30% is it extra info

OpenStudy (anonymous):

Nope, the previous answer is correct. 'Given that', so just forget what the population of the women is

OpenStudy (anonymous):

Nope, it's used to calculate part(ii)

OpenStudy (anonymous):

thanks! drawar for clarifying!

OpenStudy (anonymous):

next-ii)

OpenStudy (anonymous):

ii) P(M alcoholic)=2%

OpenStudy (anonymous):

No, it would be the correct answer if the question asks you to find the probability that the selected person is an alcoholic, GIVEN THAT the person is a man.

OpenStudy (anonymous):

how do you approach do you quantify the women in i as a chance effecting the men with the fact 2% of m2n alcoholic or I calculate like women above man independent from women?

OpenStudy (anonymous):

ii) P(M alcoholic)=100%-2%=98%

OpenStudy (anonymous):

No, use the product rule for (ii)

OpenStudy (anonymous):

ii) Product rule P(A and B)= P(A)*P(B)=2%*70%=.02*0.7=1.4%

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

thanks! can you check setup this one A poker hand (5 cards) is dealt. i Find the chance that the cards are all diamond.

OpenStudy (anonymous):

can you help me check this one: i)P(all diamond)= (13/52)*(12/51)*(11/50)*(10/49)*(9/48)

OpenStudy (anonymous):

Correct!